Determine the ratio of the electric force to the gravitational force between an electron and a proton when separated by a distance r.

1
Expert's answer
2022-02-19T06:55:20-0500

Mass of protons,mp=1.67×10-27kg

Mass of electron,me=9.10×10-31kg

Change on proton, qp=1.6×10-19c

Change on electron,qe=-1.6×10-19c


electric force:


|FE| = Kqpqe/π2

. = 9×109×1.6×10-19/π2

= 23.04×10-29/π2

Gravitational force:


|FG|=Gmpme/π2


=6.67×10-11×1.67×10-27×9.10×10-31/π2

=101.364×10-69/π2

Therefore,


|FE|/|FG|= 23.04×10-29/π2÷101.364×10-69/π2

Ans= 0.227 × 1040

Hence the ratio of electric force to the gravitational force= 0.227×1040